# Why Human Guidance Matters in Collaborative Vibe Coding

### v1

@@ Abstract
Writing code has been one of the most transformative ways for human societies to translate abstract ideas into tangible technologies. Modern AI is transforming this process by enabling experts and non-experts alike to generate code without actually writing code, but instead, through natural language instructions, or “vibe coding”. While increasingly popular, the cumulative impact of vibe coding on productivity and collaboration, as well as the role of humans in this process, remains unclear. Here, we introduce a controlled experimental framework for studying collaborative vibe coding and use it to compare human-led, AI-led, and hybrid groups. Across 16 experiments involving 604 human participants, we show that people provide uniquely effective high-level instructions for vibe coding across iterations, whereas AI-provided instructions often result in performance collapse. We further demonstrate that hybrid systems perform best when humans retain directional control (providing the instructions), while evaluation is delegated to AI.
@@
Keywords: Human-AI Collaboration; Complementarity; Vibe Coding; Large Language Models;
Figure 1: Vibe coding experimental paradigm. (A) Core procedure. An instructor views a reference image and its best SVG rendition from the previous iteration, and uses natural language instructions to guide code-generation. The Code Generator produces SVG code that can then be rendered into an image. (B) Iterated procedure. At each iteration, a selector chooses whether the current or the previous SVG image better matches the reference image. The selected SVG is then passed to an instructor, who provides vibe-coding instructions that are carried forward to the next iteration. (C) Interface of the human validation experiments. Participants rate the similarity of generated SVGs to the reference image.
Introduction
@@
Coding is a foundational skill in contemporary digital society, supporting a wide range of economic, scientific, and creative activities [18, 14]. In recent years, advances in large language models have led to rapid growth in AI-assisted programming tools, accompanied by major investment from companies [43, 17]. Systems such as Copilot, Cursor, and related tools are now widely used in real-world development workflows [42, 19]. These tools do more than speed up coding and may even act as collaborative “thought partners” [7]. They may change how people approach problems, how tasks are structured, and how decisions are made when people build.
@@
One new practice in human-AI collaborative coding is “vibe coding”. Vibe coding describes a conversational and intuition-driven style of programming wherein users provide high-level guidance and rely on iterative interaction with AI systems, rather than precise low-level control [31, 33, 10, 11, 8]. Instead of writing code, users describe goals, intentions, or desired changes, while AI systems generate and modify code accordingly. This interaction style reframes coding as a collaborative process between humans and AI, rather than a purely human activity.
@@
These developments point toward the emergence of “hybrid societies” in which humans and AI agents jointly produce complex artifacts [34, 13]. In such systems, performance depends not only on the capabilities of individual AI models, but also on how humans and AI agents coordinate over time, how responsibilities are divided, and how guidance, evaluation, and execution are distributed across agents [41, 23, 33]. Despite the importance of these questions, we still have a limited empirical understanding of how AI-assisted coding reshapes collective behavior. Most prior work focuses on improving model accuracy, interface design, or single interaction usability, rather than the iterative process of collaboration and coordination in coding [22, 1, 2].
@@
This gap persists in part because human–AI collaboration is difficult to study in the wild at scale [36]. Many studies rely on simulations of fully autonomous AI agents or offline benchmarks [40, 24], leaving open the question of whether direct human involvement – and what type of human involvement – changes outcomes in systematic ways. This raises two central questions: (i) is there a unique advantage to human–AI collaboration, rather than relying on fully automated AI pipelines?, and (ii) how should labor be effectively divided within groups that include both humans and AI?
@@
To address these questions, we introduce a controlled experimental framework for studying collaborative vibe coding. Our approach affords full causal control over who provides guidance, who executes code, who validates outputs, and how information flows across iterations. We design a task that is intuitive for non-experts, supports repeated interaction, and allows for objective validation against a known target (see Figure 1). In this paradigm, we systematically compare human- and AI-driven vibe coding, analyze sources of misalignment, test hybrid collaboration strategies, and examine how different role allocations affect collective performance.
Our approach
@@
Writing code allows people to translate arbitrary ideas into actions performed by machines. In this paper, we focus on one type of idea transfer, imagining a person wants to create a digital image. Here, provide a reference image as a stand-in for the idea people may want to communicate (Fig. 1A). The goal of the coding task is to materialize this visual idea through code. For simplicity, we use scalable vector graphics (SVG) as the coding medium. SVG provides a symbolic code that can be directly rendered into a visible image, allowing to easily judge whether the output matches the intended vision. This makes SVG well-suited for studying collaborative “vibe coding”, as participants can interpret the code by immediately observing its visual output. In this setting, a code instructor uses natural language to guide a code-generation system to produce SVG code that visually approximates the reference image. The process is iterative, emulating real-world coding practices: each instructor starts from the best result of the previous iterations and incrementally refines it to better match the target image.
Our goal is to determine under which conditions human–AI collaboration is most effective, and to assess what contribution people provide over fully automated AI coding systems.
Figure 2: Case study of human-led (human selectors and instructors) and AI-led (AI selectors and instructors) vibe coding. (A) Example progressions of the experiment for one reference image with human-led (top) and AI-led (bottom) vibe coding. (B) Examples from the last iteration of human-led (top) and AI-led (bottom) chains, more examples in Appendix Fig. 11 to 15.
Methods
Vibe coding experimental paradigm
@@ Methods
Fig. 1A illustrate our setup (see screenshots from the experiment in Appendix Fig. 8B). An instructor (vibe coder) views a reference image and its SVG rendition from the previous iteration and then uses natural language instructions to guide a code-generation system. The code-generator produces SVG code that can then be rendered as an image. We created 10 reference images by prompting GPT-5 [35] to generate images of the following animals: cat, dog, tiger, bird, elephant, penguin, shark, zebra, giraffe, panda.
@@
We also introduce a selection step to mimic realistic collaborative coding workflows. At each stage, the selector compares the current version with the previous one and can revert the changes if they do not improve the result. This ability to select and discard modifications reflects common processes of technical refinement, such as those used in version control systems (e.g., Git). Fig. 1B show the overall iterative procedure (see screenshots in Appendix Fig. 8C). At each iteration, a selector decides whether the current or the previous SVG better matches the reference image. The selected SVG is then passed to an instructor, who provides vibe-coding instructions that are passed forward.
@@
In the basic experiment, the same participant performed both the selection and instruction roles at iteration $n$. Furthermore, in the first iteration, no previous SVG is provided, and the instructor generates an image based solely on the reference image. Likewise, in the second iteration, no selector is involved, as only a single image has been created at that point.
@@
Finally, we simulate the role of independent evaluators by involving independent group of humans. These evaluators view both the reference image and the generated SVG output and assess the quality of the vibe-coding process by rating the visual similarity between the two images (Fig. 1C). Overall 120 people evaluated 4800 images (aggregated from all 16 experiments reported in this paper).
All experiments were implemented using PsyNet [12], an online framework for behavioral experiments. PsyNet naturally supports interactions between humans and AI agents within social networks [34, 26]. Appendix Fig. 16, and Appendix Fig 17 shows example of prompts provides to AI and human participants.
Participants and AI queries
Participants were recruited online via Prolific [30]. The participants were from the UK and identified English as their native language. We recruited 604 human participants overall. To simulate AI agents, we performed 4,800 API queries for code generation, and 5,327 API queries for instruction (Fig. 1A). The main AI model used was OpenAI GPT-5 [35], but for comparison we also performed additional 280 queries to Anthropic Claude Opus 4.5 [3] and 280 queries for Google Gemini 3 Pro [38]. Participants provided informed consent under the Cornell University–approved protocol IRB0148995, and were compensated at a rate of £10 per hour.
Results
Comparing human and AI roles
Figure 3: Performance of human- and AI-led vibe coding. Data points represent average rating scores across all experiments. Shaded area represents one standard error of the mean.
@@ Results
30 independent “chains” with 15 iterations were run (three repeats for the 10 reference images). We recruited 45 human participants, each of whom served as both selector and instructor across 10 chains (human-led vibe coding condition). Within each chain, the same participant first acted as the selector and then as the instructor. In a second condition (AI-led vibe coding), API calls to OpenAI’s GPT-5 model replaced human participants, while all other aspects of the experiment remained identical. In all cases, instructors and selectors used natural language to prompt a code-generation system (GPT-5).
@@
To test this quantitatively, we conducted validation experiments (Fig. 3; design described in Fig. 1C ). In the first iterations, people and AI systems attained similar ratings ($\Delta$=0.094, pp=1.000, Cohen’s dd=0.178), however in the final iteration humans significantly surpassed AI ($\Delta$=0.953, pp<0.001, Cohen’s dd=2.173). When human selectors and instructors were involved, similarity scores increased generally across iterations (indicated by a positive correlation rr=0.237, $\text{CI}_{95\%}$=[0.077, 0.387]) with highest relative improvement reaching 27.1%.
In contrast, in the AI-led baseline, similarity scores significantly decreased across iterations (indicated by a negative correlation rr=-0.351, $\text{CI}_{95\%}$=[-0.493, -0.183]). These results demonstrate a clear advantage of humans acting as selectors and instructors over a fully AI-driven coding pipeline. Moreover, while humans exhibit iterative refinement, AI systems deteriorate.
@@
As we are using a single model in the AI-led condition while involving different participants in the human-led condition, the performance difference may be attributed to the collective intelligence of human groups. To test this, we recruited another 10 participants, each independently completing 3 chains (Asocial condition; Fig. 2C) . The performance was nearly identical, and increased continuously (rr=0.289, $\text{CI}_{95\%}$=[0.130, 0.436]). These results rule out individual differences in vibe coding ability as the primary driver of performance gains and suggest that improvement arises from alignment with the task objective rather than from social interaction per se. In other words, performance gains reflect human-guided iterative refinement and are largely independent of whether improvements are produced by a single individual or by a group. To examine whether the performance difference between AI and human instructors reflects a misalignment in their object representations [37], we conducted an additional experiment in which the evaluator role was replaced by multiple GPT-5 API calls (Fig. 10A). Specifically, we asked whether AI agents would recognize that their own outputs are inferior to those produced by humans, or instead show a preference for their own creations, which would indicate a potential alignment issue.
@@
Figure 4: Comparing instruction semantics in human-led and AI-led conditions. (A) Example human- and AI-generated instructions. (B) UMAP projection of instructions in the embedding space and example word clouds of both human and AI instrubtions. (C) Validation experiment results from different AI-led experiments under instruction length limits. (D) Radar plot of seven semantic metrics: topic entropy (the diversity and unpredictability of topics), descriptive ratio (proportion of descriptive words), sentiment compound (overall emotion of the text), main IDF (how unique the vocabulary is), mean content length (average length of the answer), type token ratio (ratio of unique words to the total answer) and content ratio (proportion of content words vs. function words). For comparison, each metrics was normalized to the scale of 0 to 1.
We found a positive correlation between human and AI ratings (rr=0.348, pp<0.001; Fig. 10B) which is comparable to human and AI split halve reliability (humans: rr=0.340, AI: rr=0.448, both pp<0.001). However, AI evaluators assigned higher ratings to AI-generated outputs ($\beta$=0.298, SE=0.047, tt(895)=6.33, $\text{CI}_{95\%}$=[0.205, 0.390], pp<0.001). These findings suggest that the observed performance differences may stem from a misalignment in representations between humans and AI [15].
Figure 5: Hybrid Human-AI-Led vibe coding. (A) The schematics of human-AI hybrid vibe coding. (B) Validation rating results of human-AI hybrid vibe coding. (C) Trade-off between human proportion and the final vibe coding performance.
Divergences between human and AI instructions
@@
To further explore what drove differences in the outcomes of human and AI instructions, we next analyzed the text of the instructions. Fig. 4A shows representative examples of human and AI instructions. Human instructions are typically short and goal-directed. In contrast, AI instructions are much longer and more complex. They often contain exhaustive descriptions of visual attributes, including fine-grained details about texture, lighting, color gradients, facial features, and animal-specific anatomical properties.
@@
We next quantified semantic differences between human and AI instructions. We converted all instructions into sentence embeddings using GPT text-embedding-3-small and projected them into a two-dimensional space using UMAP ([27]; Fig. 4B). Human instructions form a single dense semantic cluster, indicating high consistency across targets and iterations. In contrast, AI instructions split into multiple target-specific clusters. This suggests that humans reuse a shared, task-oriented language across different targets, while AI treats each target animal image as a separate descriptive problem. Fig. 4B, show TF–IDF word clouds [21] demonstrating this phenomenon. Human word clouds are dominated by general action words such as “make”, “add”, “body”, and “head” which are reused across targets. AI word clouds contain many target-specific and highly descriptive terms. We then computed, across animal categories, the cosine similarity with average instruction embeddings. Human instructions showed higher similarities across categories (mean similarity M=0.841, $\text{CI}_{95\%}$=[0.828, 0.854]) compared with AI instructions (M=0.779, $\text{CI}_{95\%}$=[0.761, 0.794]). The difference was significant ($\Delta$=0.062, pp<0.001, Cohen’s dd=1.162).
@@
AI instructions were much longer than human instructions (on average instruction length for humans was M=17.723, SD=16.166, and for AI was M=755.279, SD=193.803, pp<0.001 Cohen’s $d$ = 6.479). We next tested whether instruction length alone explains the performance gap. We reran the AI-led vibe coding condition with hard limits of 10, 20, or 30 words per instruction (Fig. 4C). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, limiting instruction length did not improve AI performance, continuing to show flat or declining performance over iterations (rr=-0.202, $\text{CI}_{95\%}$=[-0.372, -0.020], r=-0.219, $\text{CI}_{95\%}$=[-0.364, -0.068], r=-0.229, $\text{CI}_{95\%}$=[-0.382, -0.068] for 10, 20 and 30 iterations), indicating that verbosity alone does not account for the observed failure. Moreover, the semantic clustering of AI instructions remained substantially tighter than that of human instructions even when controlling for instruction length. For descriptions limited to 10, 20, and 30 words, AI instructions showed significantly smaller clustering than human instructions (within category similarity means M=0.639, 0.725, 0.756, respectively; compared with human’s M=0.396, all differences pp<0.001), indicating that AI remains overly target-specific even when constrained to short descriptions.
@@
To further characterize the misalignment, we calculated seven semantic metrics for all instructions (Fig. 4D). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, AI instructions showed a higher descriptive ratio, indicating a greater proportion of descriptive words than action-oriented content. This suggests that AI focuses on detailed depiction of the target rather than on communicating effective guidance to the coding assistant. In addition, humans exhibited higher topic entropy than all AI conditions, indicating greater diversity and flexibility in how they approached improvements across iterations. This pattern suggests that humans are better at identifying what is missing or incorrect in the current SVG and at exploring alternative ways to fix it, whereas AI remains locked into narrow, over-specified descriptions.

### v2

@@ Abstract
Writing code has been one of the most transformative ways for human societies to translate abstract ideas into tangible technologies. Modern AI is changing this process by enabling experts and non-experts alike to generate code without actually writing it, instead using natural language instructions or “vibe coding”. While increasingly popular, the impact of vibe coding on productivity and collaboration, and the role of humans in this process, remains unclear. Here, we introduce a controlled experimental framework for studying collaborative vibe coding and use it to compare human-led, AI-led, and hybrid groups. Across 20 experiments involving 737 human participants, we show that people provide uniquely effective high-level instructions for vibe coding, whereas AI-provided instructions often result in performance collapse. We further demonstrate that hybrid systems perform best when humans lead by providing instructions while evaluation is delegated to AI. Although AI systems can rapidly optimize performance for specific tasks, our work highlights the importance of human guidance in shaping future hybrid societies.
@@
Keywords: Human-AI Collaboration; Complementarity; Vibe Coding; Large Language Models
1 Introduction
@@
Coding is a foundational skill in contemporary digital societies, supporting a wide range of economic, scientific, and creative activities [kelly2023starting, hu2024programming]. In recent years, advances in large language models have led to rapid growth in AI-assisted programming tools, accompanied by major investments from companies [wong2023natural, jiang2024survey]. Systems such as Copilot, Cursor, and related tools are now widely used in real-world development workflows [wermelinger2023using, kumar2025evaluating]. These tools do more than speed up coding and may even act as collaborative “thought partners” [collins2024building]. This has the potential to not only change the kind of code people write, but how they approach coding as a practice.
@@
One new practice in human-AI collaborative coding is vibe coding. Vibe coding describes a conversational and intuition-driven style of programming whereby users provide high-level guidance and rely on iterative interaction with AI systems, rather than having the traditional precise low-level control over their code [ray2025review, sapkota2025vibe, fawzy2025vibe, geng2025exploring, collins2025ai]. Instead of writing code, users describe goals, intentions, or desired changes, while AI systems generate and modify code accordingly. This interaction style reframes coding as a collaborative process between humans and AI, rather than a purely human activity.
@@
These developments point toward the emergence of “hybrid societies” in which humans and AI agents jointly produce complex artifacts [shiiku2025dynamics, hoche2025ai]. In such systems, performance depends not only on the capabilities of individual AI models, but also on how humans and AI agents coordinate over time, how responsibilities are divided, and how guidance, evaluation, and execution are distributed across agents [wang2025ai, liu2024toward, sapkota2025vibe]. Despite the importance of these questions, we still have a limited empirical understanding of how AI-assisted coding reshapes collective outcomes. Most prior work focuses on improving model accuracy, interface design, or single interaction usability, rather than the iterative process of collaboration and coordination in coding [li2025large, alam2024enhancing, amershi2019guidelines].
@@
This gap persists in part because human–AI collaboration is difficult to study in the wild at scale [sucholutsky2025using]. Many studies rely on simulations of fully autonomous AI agents or offline benchmarks [tufano2024autodev, lyu2025automatic], leaving open the question of whether direct human involvement – and what type of human involvement – changes outcomes in systematic ways. This raises two central questions: (i) is there a unique advantage to human–AI collaboration over fully automated AI pipelines?; and (ii) how should labor be effectively divided within groups that include both humans and AI?
@@
To begin to address these questions, we introduce a new controlled experimental framework for studying collaborative vibe coding. Our approach affords full causal control over who provides guidance, who executes code, who validates outputs, and how information flows across iterations. We deliberately design a task that is intuitive for non-experts, supports repeated interaction, and allows for objective validation against a known target. In this task, we systematically compare human- and AI-driven vibe coding, analyze sources of misalignment, test hybrid collaboration strategies, and examine how different role allocations affect collective performance. Our goal is to determine under which conditions human–AI collaboration is most effective, and to assess what contribution people provide over fully automated AI coding systems.
1.1 Experimental Approach
@@
Writing code allows people to translate abstract ideas into actions performed by machines. In this paper, we focus on one type of idea transfer: translating a visual idea (e.g., a mental image) into a graphical representation. In our setting, we provide a reference image as a stand-in for the idea people may want to realize (Fig. 1), and the goal of the coding task is to recreate this visual idea through code. For simplicity, we use scalable vector graphics (SVG) as the coding medium. SVG is a programming language that can be directly rendered into an image, making it possible to easily judge whether the output matches the intended target. This makes SVG well-suited for studying collaborative vibe coding, as participants can interpret the code by immediately observing its visual output. We created 10 reference images by prompting GPT-5 [singh2025openai] to generate images of the following animals: cat, dog, tiger, bird, elephant, penguin, shark, zebra, giraffe, panda.
Figure 1: Vibe coding experimental paradigm. Participants generate SVG code, which can be rendered into images. (A) Core procedure. An instructor views a reference image and its best SVG rendition from the previous iteration, and uses natural language instructions to guide code-generation. The code generator produces SVG code that is then rendered into an image. (B) Iterated procedure. At each iteration, a selector chooses whether the current or the previous SVG image better matches the reference image. The selected SVG image is then passed to an instructor, who provides vibe-coding instructions that are carried forward to the next iteration. (C) Interface of the human validation experiments. Participants rate the similarity of generated SVG images to the reference image.
@@
Fig. 1A illustrates the setup used in our experiments (see screenshots from the experiment in Appendix B). An instructor (vibe coder) views a reference image and its SVG rendition from the previous iteration and then uses natural language instructions to guide a code-generation system. The code-generator produces SVG code that can then be rendered as an image. The process is iterative, emulating real-world coding practices: each instructor starts from the best result of the previous iterations and incrementally refines it to better match the target image.
@@
We also introduce a selection step to mimic realistic collaborative coding workflows. At each stage, the selector compares the current version with the previous one and can revert the changes if they do not improve the result. This ability to select and discard modifications reflects common processes of technical refinement, such as those used in version control systems (e.g., Git). Fig. 1B shows the overall iterative procedure (see screenshots in Appendix B). At each iteration, a selector decides whether the current or the previous SVG better matches the reference image. The selected SVG is then passed to an instructor, who provides vibe-coding instructions that are passed forward.
@@
In the basic experimental configuration, the same participant performed both the selection and instruction roles at each iteration. In the first iteration, no previous SVG was provided, and the instructor generated an image solely based on the reference image. Likewise, in the second iteration, no selector was involved, as only a single image had been created at that point.
@@
Finally, we simulate the role of independent evaluators by involving an independent group of human participants. These evaluators view both the reference image and the generated SVG output and assess the quality of the vibe-coding process by rating the visual similarity between the two images (Fig. 1C, Appendix C). As an additional control, we also had AI systems perform the evaluation task. An extended Methods section is provided in Appendix A. All experiment codes and data are available at https://doi.org/10.17605/OSF.IO/BZDRV.
2 Results
2.1 Comparing human and AI roles
@@
A total of 30 independent vibe-coding “chains” with 15 iterations were run (three repeats for the 10 reference images). In one condition, we recruited 49 human participants serving as both selectors and instructors (human-led vibe coding condition). Within each chain, the same participant first acted as the selector and then as the instructor. In a second condition (AI-led vibe coding), OpenAI’s GPT-5 model performed the same task as human participants, with all other aspects of the experiment remaining identical. In all cases, instructors and selectors used natural language to prompt a code-generation system (GPT-5).
Figure 2: Example of human-led (human selectors and instructors) and AI-led (AI selectors and instructors) vibe coding. (A) Example progressions of the experiment for one reference image with human-led (top) and AI-led (bottom) vibe coding. (B) Examples from the last iteration of human-led (top) and AI-led (bottom) chains. Additional examples in Appendix E.
@@
Figure 3: Performance of human- and AI-led vibe coding. Data points represent average rating scores across all experiments. An additional asocial control whereby individual participants complete the full vibe-coding process (within-participant design) is also provided. Shaded area represents 95% confidence interval.
@@
To test this quantitatively, we recruited a separate group of human evaluators to rate the resulting images based on their similarity to the reference image (Fig. 3; design described in Fig. 1C). In the first iterations, humans and AI systems attained similar ratings ($\Delta=.06$, $p=.304$, Cohen’s $d=.08$; statistical methods described in Appendix A.3), however in the final iteration, humans significantly surpassed AI ($\Delta=1.01$, $p<.001$, $d=1.49$). When human selectors and instructors were involved, similarity scores increased generally across iterations (indicated by a positive correlation between rating scores and iterations: $r=.25$, $\text{CI}_{95\%}$=[.16, .33]) with relative improvement reaching 23.4%. In contrast, in the AI-led baseline, similarity scores significantly decreased across iterations (indicated by a negative correlation r=.23r=-.23, $\text{CI}_{95\%}$=[-.31, -.13]). These results demonstrate a clear advantage of humans acting as selectors and instructors over a fully AI-driven coding pipeline. Moreover, while humans exhibit iterative refinement, AI systems deteriorate.
@@
As we are using a single model in the AI-led condition while involving different participants in the human-led condition, the performance difference may be attributed to the collective intelligence of human groups. To test this, we recruited another 10 participants, each independently completing 3 chains (Asocial condition; Fig. 2C). The performance was nearly identical and increased continuously ($r=.12$, $\text{CI}_{95\%}$=[.03, .21]). These results rule out individual differences in vibe coding ability as the primary driver of performance gains and suggest that improvement arises from alignment with the task objective rather than from social interaction per se. In other words, performance gains reflect human-guided iterative refinement and are largely independent of whether improvements are produced by a single individual or by a group.
To examine whether the performance difference between AI and human instructors reflects a misalignment in their visual representations [sucholutsky2023getting], we conducted an additional experiment in which the evaluator role was replaced by GPT-5 (Appendix D). Specifically, we asked whether AI agents would recognize that their own outputs are inferior to those produced by humans, or instead show a preference for their own creations, which would indicate a potential alignment issue.
Overall, AI evaluators assigned higher ratings than human evaluators across all creative outputs ($\Delta=2.083$, $p<.001$, $d=2.640$). However, AI evaluators were less sensitive in discriminating between human- and AI-led stimuli (Over all iterations, human vs. AI, as evaluated by AI: $\Delta=-.06$, $p=.899$, d=.08d=-.08; human vs. AI, as evaluated by humans: $\Delta=.68$, $p<.001$, $d=.91$). Notably, in the early iterations, AI rated a higher score in AI-led vibe coding results than human-led results ($\Delta=.34$, $p<.001$, $d=.49$). This asymmetry emerged despite initially minimal baseline differences in human evaluations of human versus AI outputs, as mentioned above. Together, these findings indicate that, particularly in early iterations, AI evaluators exhibit a measurable preference for their own outputs. These findings suggest that the observed performance differences may stem from a misalignment in representations between humans and AI or from reduced overall sensitivity of AI to quality differences [huang2024characterizing].
@@
Figure 4: Comparing instruction semantics in human-led and AI-led conditions. (A) Example human- and AI-generated instructions. (B) UMAP projection of instructions in a semantic embedding space and example word clouds of both human and AI instructions. (C) Validation experiment results from different AI-led experiments under instruction length limits. (D) Radar plot of seven semantic metrics: topic entropy (the diversity and unpredictability of topics), descriptive ratio (proportion of descriptive words), sentiment compound (overall emotion of the text), main IDF (how unique the vocabulary is), mean content length (average length of the answer), type token ratio (ratio of unique words to the total answer) and content ratio (proportion of content words vs. function words). For comparison, each metrics was normalized to the scale of 0 to 1. Additional details provided in Appendix A.4.
2.2 Divergences between human and AI instruction semantics
@@
To further explore what drove differences in the outcomes of human and AI instructions, we next analyzed the text of the instructions. Fig. 4A shows representative examples of human and AI instructions. Human instructions are typically short and goal-directed. In contrast, AI instructions are much longer and more complex, often containing exhaustive descriptions of visual attributes, including fine-grained details about texture, lighting, color gradients, facial features, and animal-specific anatomical properties.
@@
To compare the semantic differences between human and AI instructions, we converted all instructions into sentence embeddings using GPT text-embedding-3-small and projected them into a two-dimensional space using UMAP ([mcinnes2018umap]; Fig. 4B). Human instructions form a single dense semantic cluster, indicating high consistency across targets and iterations. In contrast, AI instructions are split into multiple target-specific clusters. This suggests that humans reuse a shared, task-oriented language across different targets, while AI treats each target animal image as a separate descriptive problem. Fig. 4B shows TF–IDF word clouds [leskovec2020mining] demonstrating this phenomenon. Human word clouds are dominated by general action words such as “make”, “body”, and “head” which are reused across targets. AI word clouds contain more target-specific and descriptive terms, for example, using words “soft”, “slightly” frequently and using “fur” for cat and “tongue” for dog. We then computed, across animal categories, the cosine similarity with average instruction embeddings. Human instructions showed higher similarities across categories (mean similarity $M=.84$, $\text{CI}_{95\%}$=[.83, .85]) compared with AI instructions ($M=.78$, $\text{CI}_{95\%}$=[.76, .79]). The difference was significant ($\Delta=.06$, $p<.001$, $d=1.16$).
@@
AI instructions were much longer than human instructions (on average number of instruction words for humans was $M=17.72$, $SD=16.17$, and for AI was $M=755.28$, $SD=193.80$, $p<.001$, $d=6.48$). We next tested whether instruction length alone explains the performance gap. We reran the AI-led vibe coding condition with hard limits of 10, 20, or 30 words per instruction (Fig. 4C). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, limiting instruction length did not improve AI performance, continuing to show flat or declining performance over iterations (r=.20r=-.20, $\text{CI}_{95\%}$=[-0.28, -0.10], r=.15r=-.15, $\text{CI}_{95\%}$=[-.24, -.06], r=.19r=-.19, $\text{CI}_{95\%}$=[-.28, -.10] for 10, 20 and 30 iterations), indicating that verbosity alone does not account for the observed failure. Moreover, the semantic clustering of AI instructions remained substantially tighter than that of human instructions even when controlling for instruction length. For descriptions limited to 10, 20, and 30 words, AI instructions showed significantly less clustering than human instructions (within category similarity means $M=.64$, .73.73, .76.76, respectively; compared with humans $M=.40$, all differences $p<.001$), indicating that AI remains overly target-specific even when constrained to short descriptions.
@@
To further characterize the misalignment, we calculated seven semantic metrics for all instructions (Fig. 4D, Appendix A.4). After applying length limits, the mean content length of AI instructions became comparable to human instructions. However, AI instructions showed a higher descriptive ratio, indicating a greater proportion of descriptive words than action-oriented content. This suggests that AI focuses on a detailed depiction of the target rather than on communicating effective guidance to the coding assistant. In addition, humans exhibited higher topic entropy than all AI conditions, indicating greater diversity and flexibility in how they approached improvements across iterations. This pattern suggests that humans are better at identifying what is missing or incorrect in the current SVG and at exploring alternative ways to fix it, whereas AI remains locked into narrow, over-specified descriptions.
2.3 Testing hybrid human-AI instructors and selectors